The speed of a stream is 5 mph. If a boat travels 72 miles downstream in the same time that it takes to travel 36 miles upstream, what is the speed of the boat in still water?
Answer by stanbon(75887)

The speed of a stream is 5 mph. If a boat travels 72 miles downstream in the same time that it takes to travel 36 miles upstream, what is the speed of the boat in still water?
----
Downstream DATA:
distance = 72 miles ; rate = b+5 mph ; time = 72/(b+5) hrs.
----
Upstream DATA:
distance = 36 miles ; rate = b-5 mph ; time = 36/(b-5) hrs
----
Equation:
time = time
72(b+5) = 36/(b-5)
-------
2/(b+5) = 1/(b-5)
Cross-multiply to get:
2b-10 = b+5
b = 15 mph (speed of the boat in still water)
===================
Cheers,
Stan H.
===================